OpenStudy (btaylor):

A curve is traced by a point P(x,y) which moves such that its distance from the point A(-1,1) is three times its distance from the point B(2,-1). Determine the equation of the curve.

OpenStudy (btaylor):

@myininaya @Carl_Pham help? I know the answer, just not how to get it!

OpenStudy (anonymous):

Use the distance formula. \[PA = 3 PB\] \[(PA)^2 = 9 (PB)^2\] \[(x-x_A)^2 + (y-y_A)^2 = 9 \{ (x - x_B)^2 + (y - y_B)^2 \}\] \[(x+1)^2 + (y-1)^2 = 9 \{ (x - 2)^2 + (y+1)^2 \} \] Reduce from there. You'll get some kind of conic section, an ellipse I think.

OpenStudy (btaylor):

its a circle. thanx